# Recursive Entropy in Thermodynamics: Establishing the Statistical-Physics Basis of the Zentropy Approach

### v1

@@ Abstract
The recursive property of entropy is well known in the field of information theory; however, the concept is rarely used in the field of thermodynamics, despite being the field where the concept of entropy originated. This work shows that the equation for entropy used in the zentropy, which is an exact multiscale approach to thermodynamics, is a statement of the recursive property of entropy. Further, we clarify the meaning of entropy as the uncertainty arising from unconstrained degrees of freedom and separate configurational contributions from intra-configurational ones. Building on this, we derive the partition function, as used in zentropy, by maximizing entropy in its recursive form. The resulting framework is exact for a chosen level of description and enables principled coarse-graining, thereby reducing computational complexity while preserving thermodynamic consistency. These results position zentropy as a rigorous bridge between microscopic and macroscopic behavior, facilitating quantitative predictions and the study of emergent phenomena.
@@ 1 Introduction
In the present work, we set firm the mathematical foundations of the zentropy approach [Liu2024b, Liu2024c, Liu2022, Shang2023, Liu2019]. In discussions with colleagues, we have at times been met with bewilderment with regard to the entropy equation used in the zentropy approach:
@@
At first glance, there appears to be an extra term $\sum_{k}p_{k}S_{k}$ that does not belong since entropy is defined as
$S=-k_{B}\sum_{x}p_{x}\ln p_{x}.$ (2)
@@
However, we will show that Eq. (1) is equivalent to Eq. (2) if we group outcomes $x$ into $k$ groups. Furthermore, there is a long history behind Eq. (1), and it has been readily used in the fields of information theory and quantum mechanics, where it is primarily known as the recursive property of entropy, which is a hierarchical form of the chain rule of entropy. The zentropy approach can be defined as the use of the recursive property to group entropy into more useful configurations, particularly in the context of quantum mechanics and density functional theory where zentropy was introduced.
@@
The recursive property and chain rule are fundamental concepts in the fields of information theory and quantum information theory. They can be found at least as far back as 1938 [Tolman1938], and subsequently immortalized by Shannon in 1948 [Shannon1948]. However, the concept is rarely used in thermodynamics. It will be shown that restating entropy in the form of Eq. (1), enables a hierarchical coarse-graining approach to limit degrees of freedom, and greatly reduce computational complexity. The following textbooks are particularly helpful for general information on entropy and the recursive property. [Bengtsson2017, Nielsen2012, Cover2005]
@@
In the remainder of the present work, we will (i) discuss the meaning of entropy and the recursive property, (ii) show an illustrative example of the recursive property of entropy to aid intuition, (iii) provide rigorous mathematical proof of the chain rule and the recursive property, (iv) derive the partition function based on recursive entropy as used in zentropy, and (v) discuss the usefulness of zentropy as an approach for emergent phenomena and reducing computational complexity. Note that in the following sections, we set the Boltzman constant equal to unity, $k_{B}=1$, so that temperature and energy share units. We anticipate that the present work will motivate members of the thermodynamics and condensed matter community to make use of the recursive property and zentropy.
2 The meaning of entropy and the recursive property
The combined first and second law of thermodynamics at equilibrium for closed systems under hydrostatic pressure may be written
dU=TdSPdVdU=TdS-PdV (3)
@@ 2 The meaning of entropy and the recursive property
where temperature $T$ and pressure $P$ are potential quantities and entropy $S$ and volume $V$ are molar quantities. We intentionally avoid the use of intensive and extensive variables because normalized extensive variables are intensive [Hillert2007, Liu2024d, Liu2025] In many thermodynamics equations, $S$ and $V$ have symmetric places as molar quantities, and thermodynamics, per se, does not provide insight into the nature of volume and entropy. Instead, we turn to mechanics and statistical physics for answers. While volume is readily understood, entropy is more mysterious. Explaining entropy, first and foremost, we note that entropy is an anthropomorphic quantity [Jaynes1965]. It is a property of the measurement performed on the system. Entropy does not appear in Newton’s laws of motion, Maxwell’s equations, the Schrödinger equation, or Einstein’s field equations. Entropy only appears when there are degrees of freedom left unconstrained in the performance of a measurement. (While quantum mechanics contains measurement uncertainty due to wavefunction collapse, there is no uncertainty in the state of the wave function itself.) Specifically, entropy quantifies our uncertainty in the state of a system. It is important to note that this uncertainty is not from the inaccuracy of our instruments, but from unconstrained degrees of freedom of the system. When a measurement is performed on a system, usually, we do not fully constrain all degrees of freedom of the system. Instead, we reduce it to a statistical ensemble of possible configurations that comport with our measurement. For example, the canonical ensemble constrains the number of particles, volume, and temperature; it therefore includes all possible configurations compatible with these constraints regardless of other non-constrained degrees of freedom, say particle positions. All this is to explain that the entropy of the system depends on the constrained degrees of freedom. Now, what if we constrain a particular degree of freedom $k$ such that the entropy with the constraint is $S_{k}$ and the entropy without the constraint is $S$? The relation between the two is given exactly by Eq. (1).
3 Illustrative example
@@
First, let us consider an illustrative example to intuitively understand what we mean by the recursive property using probability tree diagrams as shown in Fig. 1. [Bengtsson2017ch2]
@@
In the standard approach, we consider the probability vector $\vec{p}$ with components $p_{x}$, which describes the probability of each pure state. In this way, we write the entropy as,
$S(\vec{p})=-\sum_{x}p_{x}\ln p_{x}$ (4)
Now, instead of concerning ourselves with every possible outcome, let us group outcomes, and consider the probability distribution for each group using the probability vector $\vec{q}$ with components
$q_{k}=\sum_{x\in G_{k}}p_{x}$ (5)
where $G_{k}$ is the set of all the pure states for the $k$th group and each group is distinct so that $\{G_{k}\}$ forms a partition of the pure states (assumed here for intuitive ease, the proofs of Section 4 do not have this requirement). In this way, the entropy of the groups is
$S(\vec{q})=-\sum_{k}q_{k}\ln q_{k}$ (6)
This is often referred to as the configurational entropy in the literature [AdamGibbs1965, MiracleSenkov2017]. Of course, some entropy remains after choosing a particular group, and it is obvious that $S(\vec{q})\neq S(\vec{p})$. The entropy that remains after choosing a group is
$S_{k}\left(\frac{\vec{p}}{q_{k}}\right)=-\sum_{x\in G_{k}}\frac{p_{x}}{q_{k}}\ln\frac{p_{x}}{q_{k}}$ (7)
and represents the entropy that remains after constraining a particular degree of freedom $k$. We call this the intra-configurational entropy. The total entropy $S(\vec{p})$ can then be written as the configurational entropy and the weighted sum of the intra-configurational entropies,
$S(\vec{p})=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)\\$ (8)
This can be shown using the fact that $p_{x}=q_{k}\cdot\frac{p_{x}}{q_{k}}$ and the logarithmic product rule $\ln{ab}=\ln a+\ln b$. Starting from the entropy of the pure states,
@@
$=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right)$ (15)
In this illustration, we partitioned pure-state configurations into groups. For this reason, the recursive property is sometimes called the grouping property. The justification for the recursive name is that we could just as easily group mixed-state configurations—group the groups, so to speak. Then the total entropy would be calculated recursively from subordinate scales.
4 Derivation of recursive entropy
To derive the entropy equation used in the zentropy approach, Eq. (1), we first use several definitions and theorems with the random variable formalization, as it facilitates conceptual progression. First, we start by defining the Shannon entropy [Cover2005]. Let $X$ be a random variable with outcomes $x$, alphabet $\mathcal{X}$, and probability mass function $p(x)$. For short, $p(x)=\Pr\{X=x\},x\in\mathcal{X}$.
@@ 4 Derivation of recursive entropy
$S(X)=-\sum_{x\in\mathcal{X}}p(x)\ln p(x)$ (16)
or equivalently
$S(p)=-\sum_{x}p_{x}\ln p_{x}$ (17)
where we define $0\ln 0\equiv 0$, a convention justified by noting that $\lim_{x\to 0}x\ln x=0$ and the intuition that a state $x$ with $p_{x}=0$ should not contribute to the entropy.
Next, we extend the definition of Shannon entropy to a pair of discrete random variables [Cover2005].
@@
The joint entropy of a pair of discrete random variables $(X,Y)$ with a joint distribution $p(x,y)$ is defined by
@@
Additionally, we consider the entropy of $Y$ conditional on knowing $X$. The entropy of $Y$ given $X$ is naturally the entropies of the conditional distributions averaged over the conditioning random variable [Cover2005].
@@
If the random pair $(X,Y)$ has a joint distribution $p(x,y)$ the conditional entropy is defined as
@@
$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(y\mid x).$ (21)
We now present the first version of the chain rule, which states that the entropy of a pair of random variables can be expressed as the entropy of one variable plus the conditional entropy of the other [Cover2005].
@@
Starting with Eq. (18) for the definition of joint entropy,
@@
$=S(X)+S(Y\mid X)\qed$ (27)
@@
For the second version of the chain rule, we generalize to $n$ random variables. By repeatedly applying the two-variable chain rule, the entropy of $n$ random variables can be expressed as the sum of conditional entropies by induction [Cover2005].
@@
$S(X_{1},X_{2},...,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})$ (28)
@@
$S(X_{1},X_{2},...X_{n})$ $=S(X_{1})+S(X_{2}\mid X_{1})+...+S(X_{n}\mid X_{n-1},...X_{1})$ (33)
$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})\qed$ (34)
Thus far, we have naturally treated each random variable on an equal level in accordance with the entropy being a symmetric function of the random variables. However, it is sometimes useful to consider one (or more) random variables hierarchically. That is, we wish to consider the entropy between configurations and the entropy within those configurations. To do this, we first define the entropy within a given configuration $k$:
@@
$S_{k}$ $=S(X_{1},X_{2},...X_{n}\mid K=k)$ (35)
$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K=k)$ (36)
Let $K$ be a random variable with probability distribution $p(K)$, where $\sum_{k}p_{k}=1$. We are now ready to prove our third version of the chain rule for the case of multiscale entropy.
@@
$S(K,X_{1},X_{2},...,X_{n})=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}$ (37)
@@
First we choose to write the chain rule in the following form using the definition of conditional entropy from equation 19.
$S(X_{1},X_{2},...,X_{n})$ $=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1})$ (38)
$S(K,X_{1},X_{2},...,X_{n})$ $=S(K)+\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K)$ (39)
$=S(K)+\sum_{k\in\mathcal{K}}p(k)\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},...,X_{1},K=k)$ (40)
$=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}\qed$ (41)
@@
Note that in this form, there is a conspicuous absence of any reference to any of the $X_{i}$ on the right-hand side as they are all subsumed into $S_{k}$. Hence, the recursive property can be regarded as the hierarchical form of the chain rule.
@@
The derivations presented in this section establish the mathematical foundation of the recursive property and show that the recursive property is not only consistent with the definition of entropy, but also essential. In fact, the recursive property is often taken as an axiom to single out the Shannon entropy as the natural choice of equations [khinchin1953concept, faddeev1956concept, Aczel1974] for entropy. The recursive property, as a hierachical form of the chain rule, reveal that the total entropy can be decomposed into a configuration component $\sum_{k}p_{k}\ln p_{k}$ and weighted sum of the intra-configurational entropies $\sum_{k}p_{k}S_{k}$. The recursive property represents a generalization of the standard entropy definition that reduces to the familiar form when configurations are chosen to be quantum pure states. That is, if $k$ specifies all degrees of freedom any other information ($X_{i}$) willbe redundant and the intra-configurational entropy is zero for all $k$, $\forall k\in\mathcal{K},\;S_{k}=0$. This flexibility in choosing the scale of description provides a powerful multiscale framework known as Zentropy for addressing systems on multiple scales where a natural partitioning of configurations exists, enabling efficient coarse-graining of degrees of freedom while maintaining thermodynamic consistency.
5 Derivation of the partition function
Definition 5.
The Helmholtz energy is defined as
F=ETSF=E-TS (42)

### v2

@@ Abstract
The recursive property of entropy is well known in the field of information theory; however, the concept is rarely used in the field of thermodynamics, despite being the field where the concept of entropy originated. This work shows that the equation for entropy used in the zentropy approach, which is an exact multiscale approach to thermodynamics, is a statement of the recursive property of entropy. Further, we clarify the meaning of entropy as the uncertainty arising from unconstrained degrees of freedom and separate configurational and intra-configurational contributions. Building on this, we derive the Helmholtz energy and the partition function, as used in zentropy. The resulting framework is exact for a chosen level of description and enables principled coarse-graining, thereby reducing computational complexity while preserving thermodynamic consistency. These results position zentropy as a rigorous bridge between microscopic and macroscopic behavior, facilitating quantitative predictions and the study of emergent phenomena.
@@ 1 Introduction
The Gibbs (Shannon) entropy is a function of the probabilities of all possible microstates of a system and quantifies the uncertainty in the state of the system,
$S=-k_{B}\sum_{x}p_{x}\ln p_{x}.$ (1)
From this definition, there is no prescription for the ordering of the microstates. The only requirement is that the summation be performed over the entire set. This leaves us with considerable freedom in our description of the system. In fact, as is done in the zentropy approach, the summation may even be performed hierarchically, with some adjustments. The zentropy approach [Liu2019, Shang2023, Hew2024, Hong2025] exploits this choice in description to facilitate hierarchical coarse-graining, and the approach was first introduced as a method of “multi-scale entropy” [Liu2019]. Later, the method adopted the name zentropy [Liu2022], where “Z” refers to the partition function $Z$, itself derived from the German word Zustandssumme, meaning “sum over states”. The term zentropy emphasizes that this approach sums over states with entropy.
[…]
In the present work, we set firm and elaborate on the mathematical foundations of the zentropy approach, where entropy is given by
@@
At first glance, there appears to be an extra term $\sum_{k}p_{k}S_{k}$. However, we will show that Eq. (2) is equivalent to Eq. (1) if we group outcomes $x$ into groups $k$. Furthermore, there is a long history behind Eq. (2), where it is primarily known as the recursive property of entropy, which is a hierarchical form of the chain rule of entropy. This property of entropy has been widely used in the fields of information theory and quantum mechanics. The zentropy approach can be defined as the use of the recursive property to group microstates into more useful coarse-grained configurations, particularly in the context of quantum mechanics and density functional theory, where zentropy was introduced.
@@
The recursive property and chain rule are fundamental concepts in the fields of information theory and quantum information theory. They can be found at least as far back as 1938 [Tolman1938, Chap. IV, p. 176], and subsequently immortalized by Shannon in 1948 [Shannon1948]. However, the concept is rarely used in thermodynamics. It will be shown that restating entropy in the form of Eq. (2), enables a hierarchical coarse-graining approach to limit degrees of freedom and greatly reduce computational complexity. The following textbooks are particularly helpful for general information on entropy and the recursive property [Bengtsson2017, Nielsen2012, Cover2005].
@@
In the remainder of the present work, we will (i) discuss the meaning of entropy and the recursive property, (ii) show an illustrative example of the recursive property of entropy to aid intuition, (iii) provide rigorous mathematical proof of the chain rule and the recursive property, (iv) derive the partition function based on recursive entropy as used in zentropy, and (v) discuss the usefulness of zentropy as an approach for emergent phenomena and reducing computational complexity. Note that in the following sections, we set the Boltzmann constant equal to unity, $k_{B}=1$, so that temperature and energy share the same units. We anticipate that the present work will motivate members of the thermodynamics and condensed matter communities to make use of the recursive property and zentropy.
2 The Meaning of Entropy: Constrained and Unconstrained Degrees of Freedom
For a simple compressible closed system at thermodynamic equilibrium, the fundamental differential relation may be written as
dU=TdSPdV.dU=TdS-PdV. (3)
In this form, $S$ and $V$ appear symmetrically as conjugate variables to $T$ and $P$, respectively, and thermodynamics per se does not provide insight into the nature of volume and entropy. Instead, we must rely on mechanics and statistical physics for answers. While volume is readily understood, entropy is more mysterious.
@@ 2 The Meaning of Entropy: Constrained and Unconstrained Degrees of Freedom
To explain entropy, first and foremost, we note that entropy is an anthropomorphic quantity [Jaynes1965]. It is a property of the measurement performed on the system. Entropy does not appear in Newton’s laws of motion, Maxwell’s equations, the Schrödinger equation, or Einstein’s field equations. Instead, it only appears when there are degrees of freedom left unconstrained when performing a measurement. (While quantum mechanics contains measurement uncertainty due to wavefunction collapse, there is no uncertainty in the state of the wavefunction itself.) Specifically, entropy quantifies our uncertainty in the state of a system. It is important to note that this uncertainty is not due to instrument inaccuracy but to the unconstrained degrees of freedom of the system. When a measurement is performed on a system, we usually do not fully constrain all degrees of freedom of the system. Instead, we reduce it to a statistical ensemble of possible microstates that comport with our measurement.
For example, the canonical ensemble constrains the number of particles, volume, and temperature; it therefore includes all possible microstates compatible with these constraints, regardless of other unconstrained degrees of freedom, such as particle positions. For clarity, “degrees of freedom” here refers to the microscopic degrees of freedom, not to the macroscopic state variable degrees of freedom of Gibbs’ phase rule. All this is to explain that the entropy of the system also depends on the constrained degrees of freedom. Now, what if we constrain a particular degree of freedom $k$ such that the entropy with the constraint is $S_{k}$ and the entropy without the constraint is $S$? The relation between the two is given exactly by Eq. (2).
3 Intuitive Illustration of Recursive Entropy: Partitioned Case
@@
To intuitively understand what we mean by the recursive property, let us consider an illustrative example using probability tree diagrams [Bengtsson2017ch2] as shown in Figure 1.
@@
In the ungrouped scenario (Figure 1a), we consider the probability vector $\vec{p}$ with components $p_{x}$, which describes the probability of each microstate. In this way, we write the entropy as,
$S(\vec{p})=-\sum_{x}p_{x}\ln p_{x}.$ (4)
Now, instead of concerning ourselves with every possible outcome, let us group outcomes (Figure 1b), and consider the probability distribution for each group or partially-constrained configuration using the probability vector $\vec{q}$ with components $q_{k}$ defined by
$q_{k}=\sum_{x\in G_{k}}p_{x},$ (5)
where $G_{k}$ is the set of all fully-constrained configurations for the $k$th group and each group is distinct so that $\{G_{k}\}$ forms a partition of the fully-constrained configurations (assumed here for intuition; the proofs in Section 4 do not require strict partitions and hold in the more general case). In this way, the entropy between the groups is
$S(\vec{q})=-\sum_{k}q_{k}\ln q_{k}.$ (6)
This is often referred to as the configurational entropy in the literature [AdamGibbs1965, MiracleSenkov2017, Hong2025, Shang2023]. Of course, some entropy remains after choosing a particular group, and it is obvious that $S(\vec{q})\neq S(\vec{p})$. The entropy that remains after choosing a group is
$S_{k}\left(\vec{p}_{|k}\right)=-\sum_{x\in G_{k}}\frac{p_{x}}{q_{k}}\ln\frac{p_{x}}{q_{k}},$ (7)
where $\vec{p}_{|k}$ is the probability distribution after choosing a particular group $k$ and has components $p_{x|k}=p_{x}/q_{k}$. $S_{k}$ represents the entropy that remains after constraining a particular degrees of freedom. We call this the intra-configurational entropy. The total entropy $S(\vec{p})$ can then be written as the configurational entropy plus the weighted sum of the intra-configurational entropies
$S(\vec{p})=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right).$ (8)
This can be shown using the fact that $p_{x}=q_{k}\cdot\frac{p_{x}}{q_{k}}$ and the logarithmic product rule $\ln{ab}=\ln a+\ln b$. Starting from the entropy of the fully-constrained configurations
@@
$=S(\vec{q})+\sum_{k}q_{k}S_{k}\left(\frac{\vec{p}}{q_{k}}\right).$ (15)
In this illustration, we partitioned fully-constrained configurations into groups. For this reason, the recursive property is sometimes called the grouping property. The term recursive is justified because we could just as easily group the partially-constrained configurations—group the groups, so to speak.
Some terminological clarification is required here. In statistical mechanics, the terms microstate and macrostate are introduced to clarify the meaning of state. In thermodynamics, the state of a system has a property called entropy. However, the fully specified microscopic description of the same system is purely mechanical, and there are no states with entropy. Thus, the macrostate refers to the system as it appears macroscopically, with many degrees of freedom left unconstrained, while the microstate refers to the system fully specified, with all degrees of freedom constrained.
This distinction becomes more subtle in the zentropy approach, in which the system has partially-constrained configurations. In this context, configurations may be viewed two ways: they can be microstates that comprise a macrostate or super-configuration, or they can be macrostates with their own microstates or sub-configurations. To avoid ambiguity, we use the term configuration to refer generally to anything that corresponds to a particular group $k$ of Eq. (2) (or Eq. (8) of this section). Further, a configuration may be called a microstate when treated as a sub-configuration and a macrostate when treated as a super-configuration.
4 Derivation of Recursive Entropy: General Case
While the previous section provided the intuition and showed the particular case where the fully-constrained configurations are partitioned, in this section, we provide a general proof of the recursive property that does not rely on forming a partition. To do this, we first establish several definitions and theorems, largely following the approach and notation of Cover and Thomas [Cover2005], starting with the definition of Shannon entropy. Let $X$ be a random variable with outcomes $x$, alphabet $\mathcal{X}$, and probability mass function $p(x)$. For short, $p(x)=\Pr\{X=x\},x\in\mathcal{X}$.
@@ 4 Derivation of Recursive Entropy: General Case
$S(X)=-\sum_{x\in\mathcal{X}}p(x)\ln p(x),$ (16)
where we define $0\ln 0\equiv 0$, a convention justified by noting that $\lim_{x\to 0}x\ln x=0$ and that the intuition that a microstate $x$ with $p(x)=0$ should not contribute to the entropy.
Next, we extend the definition of Shannon entropy to a pair of discrete random variables. Let $Y$ also be a random variable with alphabet $\mathcal{Y}$ and probability mass function $p(y)=\Pr\{Y=y\}$, and let $p(x,y)$ denote the joint probability mass function of $(X,Y)$.
@@
The joint entropy of a pair of discrete random variables $(X,Y)$ with a joint probability mass function $p(x,y)$ is defined by
@@
Additionally, we consider the entropy of $Y$ conditional on knowing $X$. This entropy is naturally the entropies of the conditional distributions averaged over the conditioning random variable.
@@
If the random pair $(X,Y)$ has a joint probability mass function $p(x,y)$ the conditional entropy is defined as
@@
$=-\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\ln p(y\mid x),$ (20)
where the last equality follows from the definition of the joint probability mass function, $p(x,y)=p(x)p(y\mid x)$.
We now present the first version of the chain rule, which states that the entropy of a pair of random variables can be expressed as the entropy of one variable plus the conditional entropy of the other.
@@
Starting from the definition of joint entropy (Eq. (17)),
@@
$=S(X)+S(Y\mid X).\qed$ (26)
In the fourth line of the derivation, we used the marginal probability relation $p(x)=\sum_{y\in\mathcal{Y}}p(x,y)$.
@@
For the second version of the chain rule, we generalize to $n$ random variables through induction by repeatedly applying the two-variable chain rule.
@@
The entropy of $n$ random variables can be expressed as the sum of conditional entropies
$S(X_{1},X_{2},\ldots,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1}).$ (27)
@@
$S(X_{1},X_{2},\ldots X_{n})$ $=S(X_{1})+S(X_{2}\mid X_{1})+\ldots+S(X_{n}\mid X_{n-1},\ldots,X_{1})$ (32)
$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1})\qed$ (33)
Thus far, we have treated each random variable on the same hierarchy, and it is natural to do so as the Shannon entropy is symmetric under permutation of the arguments. However, it is sometimes useful to consider one or more random variables hierarchically. That is, to treat some random variables as superior to others, and examine the entropy that remains when their outcomes are fixed. To do this, let $K$ be a random variable with probability mass function $p(k)$, where $\sum_{k\in\mathcal{K}}p(k)=1$.
@@
The entropy within a given configuration $k$, called the intra-configurational entropy is defined as
$S_{k}$ $=S(X_{1},X_{2},\ldots,X_{n}\mid K=k)$ (34)
$=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1},K=k).$ (35)
In this way, we can treat $K$ as superior to the other random variables $X$, and all the fully-constrained $X$-configurations are grouped by the outcomes $k$. However, note that the groups need not form a partition of the fully-constrained configurations, contrasting with Section 3, since different $k$ values may assign nonzero probability to the same $X$-configuration.
We are now ready to prove our third version of the chain rule for the case of multiscale entropy.
@@
The entropy of $n+1$ random variables may be expressed recursively as
$S(K,X_{1},X_{2},\ldots,X_{n})=-\sum_{k\in\mathcal{K}}p(k)\ln p(k)+\sum_{k\in\mathcal{K}}p(k)S_{k}.$ (36)
For convenience, we may define $p_{k}=p(k)$ and write the sum over $k\in\mathcal{K}$ simply as $\sum_{k}$:
$\boxed{S(K,X_{1},X_{2},\ldots,X_{n})=-\sum_{k}p_{k}\ln p_{k}+\sum_{k}p_{k}S_{k}.}$ (37)
@@
We begin with the chain rule:
$S(X_{1},X_{2},\ldots,X_{n})=\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1}).$ (38)
Then, we introduce $K$ and use the chain rule to separate the entropy of $K$ from the remaining variables. We then apply the definitions of conditional entropy (Eq. (18)) and intra-configurational entropy $S_{k}$ (Eq. (35)):
$S(K,X_{1},X_{2},\ldots,X_{n})$ $=S(K)+\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1},K)$ (39)
$=S(K)+\sum_{k\in\mathcal{K}}p(k)\sum_{i=1}^{n}S(X_{i}\mid X_{i-1},\ldots,X_{1},K=k)$ (40)
$=-\sum_{k\in\mathcal{K}}p(k)\ln p(k)+\sum_{k\in\mathcal{K}}p(k)S_{k}.\qed$ (41)
@@
Note that in this form, there is a conspicuous absence of any reference to any of the $X_{i}$ on the right-hand side of the last equality, as they are all subsumed into $S_{k}$. Hence, the recursive property can be regarded as the hierarchical form of the chain rule.
The derivations in this section establish the mathematical foundation of the zentropy and demonstrate the consistency of the recursive property with the definition of Shannon entropy. This property is fundamental and often used as an axiom to uniquely determine the Shannon entropy as the most natural measure of uncertainty [khinchin1953concept, faddeev1956concept, Aczel1974]. The recursive property, as a hierarchical form of the chain rule, reveals that the total entropy can be decomposed into configurational and intra-configurational terms:
$S=S_{\text{conf}}+\langle S_{k}\rangle,\\$ (42)
where the configurational entropy is the entropy on the preferred level of hierarchy defined by $K$ and the intra-configurational term is the expectation value of the intra-configurational entropies:
$S_{\text{conf}}$ $=-\sum_{k\in\mathcal{K}}p(k)\ln p(k)$ (43)
$\langle S_{k}\rangle$ $=\sum_{k\in\mathcal{K}}p(k)S_{k}.$ (44)
@@
The recursive property provides a hierarchical decomposition of the standard entropy definition that reduces to the familiar form when configurations are fully-constrained with no microscopic degrees of freedom. That is, if $k$ specifies all degrees of freedom, any other information ($X_{i}$) will be redundant, and the intra-configurational entropy is zero for all $k$ (that is, $\forall k\in\mathcal{K},\;S_{k}=0$). This flexibility in choosing the scale of description provides a powerful multiscale framework known as zentropy for addressing systems on multiple scales, enabling efficient coarse-graining of degrees of freedom while maintaining thermodynamic consistency.
5 Helmholtz energy and the Partition Function
Since the Helmholtz energy and the partition function are defined downstream from entropy, we have to take into account intra-configurational entropy when evaluating these quantities for partially-constrained configurations. Since the partially-constrained configurations contain both energy and entropy, we can think of the partially-constrained configurations as sub-ensembles. In fact, in this section, we show that the Helmholtz energy and partition function are evaluated in terms of an intrinsic Helmholtz energy $F_{k}$ for the partially-constrained case instead of energy $E_{x}$ as in the fully-constrained case.
From thermodynamics, the Helmholtz energy is defined as
F=UTS,F=U-TS, (45)